OpenStudy (anonymous):

Determine the equation of the line, in standard form, that will get you from the Launch Area to Point A,

OpenStudy (anonymous):

Launch Area:___(1, 2)___ Point A:___(0, 3)___ Point B:___(-3, 0)___ Point C:___(-1, -4)___

OpenStudy (anonymous):

I just need to know how to determine the equation. How to set up the equation, what to fill in, how to make it standard form and what not.

OpenStudy (anonymous):

@dumbcow @ganeshie8

OpenStudy (anonymous):

Do you know how to determine slope for a line between 2 points?

OpenStudy (anonymous):

y2-y1 / x2-x1 that'd mean that it'd be 3 - 2/0-1 my slope would be 1/-1

OpenStudy (anonymous):

Or just -1 correct.

OpenStudy (anonymous):

so how would I make that into standard form? standard form is A(x) + B(y) = C right? so... what would I do?

OpenStudy (anonymous):

Take one of the points in your line...the question ask for the form from Launch Pad to pint A. So you can use either coordinate. Let's use the launch pad (1,2). Using these pints we need to determine our equation in slope intercept form first.\[(y-y ^{1})=slope(x-x ^{1})\]

OpenStudy (anonymous):

Plug in your values (1,2) for y(1) and x(1) and rewrite the equation of the line.

OpenStudy (anonymous):

okay so it would look like (y - 2) = -1 (x - 1)?

OpenStudy (anonymous):

Yes..but take it a step further so you end up with y=mx+b

OpenStudy (anonymous):

y = -1(1) + 2? so it'd just be y= 1?

OpenStudy (anonymous):

Simplify first and we get y-2=-x+1 as I distributed( -1) on the right side. Now put in standard form which is Ax+By=C. by getting the x and y values to one side of the equation

OpenStudy (anonymous):

I'm kinda slow at this so please bare with me as I make a fool of myself for doing it wrong.

OpenStudy (anonymous):

okay so -2 would be added to cancel? that'd leave us with y = x+1+(-2)?

OpenStudy (anonymous):

Your doing great but I must leave..You should end up with x+y=3. This is the standard form.

OpenStudy (anonymous):

thank you for your help for the time c:

OpenStudy (anonymous):

@nelsonjedi could you come back and help me with this some more? Dx